OpenStudy (anonymous):

solve f(x)+14/x-4; x=4 Staet whether the function is continuous at the indicated point. If not why?

OpenStudy (anonymous):

solve f(x)=14/x-4; x=4 Staet whether the function is continuous at the indicated point. If not why?

OpenStudy (anonymous):

olve f(x)=14/x-4; x=4 State whether the function is continuous at the indicated point. If not why?

OpenStudy (anonymous):

if x=4, then $$ \frac{14}{x-4)}=14/0 $ so is not continuous. $$\text {Solve } f(x)=\frac{14}{x-4} $$ not sure what you mean, as f(x) is a function, hope this helps.

OpenStudy (anonymous):

$$ f(x)=\frac{14}{x-4}: x=4 =\frac{14}{0}$$ so is not continuous - hope thant's clearer

OpenStudy (anonymous):

Thanks. I'm assuming it is not contiunous; f(4) does not exist and lim of x as it goes to 4 f(x) does not exist as well.

OpenStudy (anonymous):

A function f(x) is continuous at a c if the following 3 conditions are met <definition of continuity at a point> 1) f(c) is defined 2) lim_(x->c) f(x) Exists 3) lim_(x->c) f(x) = f(c) Although we only had to look at the first rule, since f(x) is not defined at c = 4, it is not continuous at that point. Thats just part of the definition of Continuity at a Point.